OpenStudy (anonymous):

Solve 3|2x + 2| − 7 = 5 for x. Show your work

OpenStudy (anonymous):

2x + 2= +(5+7)/3 2x + 2= -(5+7)/3 solve this 2 equations for x

OpenStudy (hoblos):

3|2x + 2| − 7 = 5 add 7 for both sides 3|2x + 2| = 12 divide both sides by 3 |2x + 2| = 4 if |a| = b then a= +or- b thus 2x +2 = 4 or 2x+2 =-4 2x = 2 2x = -6 x=1 x=-3

OpenStudy (anonymous):

ok thanks